OpenStudy (anonymous):

A line passes through (1, -4) and is parallel to the line 3x-8y+1=0. Find its equation

OpenStudy (anonymous):

@pielover123

OpenStudy (igreen):

In what form?

OpenStudy (igreen):

Slope-intercept?

OpenStudy (igreen):

Standard? Point-slope?

OpenStudy (anonymous):

sorry i took so long

OpenStudy (anonymous):

any line parallel will have the equation 3x +8y +c you have to find c put in (1,-4) and solve to get c

OpenStudy (igreen):

3x - 8y + 1 = 0 Convert to slope-intercept form: Subtract 1 to both sides: 3x - 8y = -1 Subtract 3x to both sides: -8y = -3x - 1 Divide -8 to both sides: y = 3/8x + 1/8 Okay, what's the slope here? @~Gelmhar

OpenStudy (anonymous):

3/8

OpenStudy (anonymous):

sorry -8x typo

OpenStudy (anonymous):

got it :D, can we proceed to another question :D

OpenStudy (igreen):

Correct, and parallel lines have the SAME slope..so the slope of our new line will have a slope of 3/8 also. Now let's plug in (1, -4) and the slope of 3/8 into point-slope form. \(y - y_1 = m(x - x_1)\) Where \(y_1\) is the y-value of the point, \(x_1\) is the x-value of the point, and \(m\) is the slope. Can you plug in the point and the slope?

OpenStudy (igreen):

Oh, what did you come up with?

OpenStudy (anonymous):

im still solving it @iGreen

OpenStudy (igreen):

Okay, tell me what you come up with.

OpenStudy (anonymous):

-3x +8y+35=0

OpenStudy (anonymous):

or 3x-8y-35=0

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

That's incorrect..

OpenStudy (igreen):

Do you have options?

OpenStudy (anonymous):

how did you know that is incorrect?

OpenStudy (igreen):

It isn't the same as the original line.

OpenStudy (anonymous):

I get 3x-8y-35 also

OpenStudy (igreen):

Oh, wait, hold on.

OpenStudy (igreen):

That's correct, sorry..I just thought it had to be the same as that line..lol. My bad.

OpenStudy (anonymous):

i substitute it correctly :D

OpenStudy (igreen):

Nice work! \(\Huge\color{lime} \checkmark\)

OpenStudy (anonymous):

can you help me again igreen :D

OpenStudy (igreen):

Sure thing!